Please help solve for x in the following problem, I couldn't figure it out just by myself, help is appreciated
log(6x-1)=1+log(9x-6)

Answer by scott8148(6628)

log(6x-1) = log(10) + log(9x-6)

6x - 1 = (10)(9x - 6) ___ 6x - 1 = 90x - 60